Calculate the persent by mass of each element in the following:.
Calcium chloride CaCl2
Sodium hydrocarbonate NaHCo3
Ammoniom sulphat (NH4)2 SO4
Sodium thiosulphate Na2 S2 SO3
1
Expert's answer
2012-08-09T08:24:38-0400
mass percent of an element in a compound = mass of element in 1 mole of compound/mass
of 1 mole of compound x 100%
[yeah it doesn't make a whole lot of sense]

calcium chloride(CaCl2)--molar mass = (40.08) + 2(35.45) = 40.08 + 70.90 = 110.98
this is the number that goes on the bottom.


%Ca = 1 mole of Ca in molecule x 40.08 g/mole(this is the atomic mass of calcium) /
110.98 g/mole x 100% = 0.36 x 100% = 36%

%Cl = 2 moles of Cl in molecule x 35.45 g/mole(atomic mass of chlorine) / 110.98
g/mole x 100% = 0.64 x 100% = 64%

36% Calcium, 64%
Chlorine

NaHCO3

Mr=22+1+12+3*16=83

Na=22/83=0.265 26.5%
H=1/83=0.012 1.2%
C=12/83=0.145 14.5%
O=3*16/83=0.578 57.8%

(NH4)2SO4:

Element | Percentage | Number of atoms | Mass of atom
______________________________________…
Nitrogen |... 21.20%... |............2............ | 14.00674
Hydrogen |....6.10%....|............8............… 1.00794
Sulfur |.......24.27&......|...........1.......…
32.066

Na2S2

Na = 29%
S = 40.5%
O = 30%